OpenStudy (anonymous):

What Are The Solutions of The System? y = x^2 + 3x - 4 y = 2x + 2 A. (–3, 6) and (2, –4) B. (–3, –4) and (2, 6) C. (–3, –4) and (–2, –2) D. no solution I Think It's B Is That Correct?

OpenStudy (lgbasallote):

let's check B coordinate 1: (-3, -4) equation 1: -4 = (-3)^2 + 3(-3) - 4 -4 = 9 - 9 - 4 -4 = -4 equation 2: -4 = 2(-3)+ 2 -4 = -6 + 2 -4 = -4 coordiante 2:(2,6) eq 1: 6 = 2^2 + 3(2) - 4 6 = 4 + 6 - 4 6 = 6 eq 2: 6 = 2(2) + 2 6 = 4 + 2 6 = 6 THEREFORE B IS RIGHT

OpenStudy (anonymous):

That's My Exact Work I Just Needed To Make Sure Because My Brother Was Saying I Was Incorrect :D

OpenStudy (lgbasallote):

well you're right so dont worry haha